### v1

# UQLM: A Python Package for Uncertainty Quantification in Large Language Models

Abstract
Hallucinations, defined as instances where Large Language Models (LLMs) generate false or misleading content, pose a significant challenge that impacts the safety and trust of downstream applications. We introduce uqlm, a Python package for LLM hallucination detection using state-of-the-art uncertainty quantification (UQ) techniques. This toolkit offers a suite of UQ-based scorers that compute response-level confidence scores ranging from 0 to 1. This library provides an off-the-shelf solution for UQ-based hallucination detection that can be easily integrated to enhance the reliability of LLM outputs. [^1] Note: This paper is not intended to reflect the specific work, practices, or opinions of the author-affiliated company. Any resemblance to actual practices is coincidental. All opinions are the authors’ own.
Keywords: large language model, uncertainty quantification, hallucination detection, Python, AI safety
1 Introduction
Large language models (LLMs) have revolutionized the field of natural language processing, but their tendency to generate false or misleading content, known as hallucinations, significantly compromises safety and trust. LLM hallucinations are especially problematic because they often appear plausible, making them difficult to detect and posing serious risks in high-stakes domains such as healthcare, legal, and financial applications. As LLMs are increasingly deployed in real-world settings, monitoring and detecting hallucinations becomes crucial.
Traditional evaluation methods involve ‘grading’ LLM responses by comparing model output to human-authored ground-truth texts, an approach offered by toolkits such as Evals (OpenAI, 2024) and G-Eval (Liu et al., 2023). While effective during pre-deployment testing, these methods are limited in practice since users typically lack access to ground-truth data at generation time. This shortcoming motivates the need for generation-time hallucination detection methods.
Existing solutions to this problem include source-comparison methods, internet-based grounding, and uncertainty quantification (UQ) methods. Toolkits that offer source-comparison scorers, such as Ragas (Es et al., 2023), Phoenix (Arize AI, 2025), DeepEval (Ip and Vongthongsri, 2025), and others (Hu et al., 2024; UpTrain AI Team, 2024; Zha et al., 2023; Asai et al., 2023) evaluate the consistency between generated content and input prompts. However, these methods can mistakenly validate responses that merely mimic prompt phrasing without ensuring factual accuracy. Toolkits that leverage Internet searches for fact-checking, such as FacTool (Chern et al., 2023), introduce delays and risk incorporating erroneous online information, failing to address the inherent uncertainty in model outputs. Lastly, although numerous UQ techniques have been proposed in the literature, their adoption in user-friendly, comprehensive toolkits remains limited. For example, while SelfCheckGPT (Manakul et al., 2023) incorporates some UQ scorers, its set of techniques is narrow and does not integrate generation with evaluation, thus creating barriers for practitioners outside specialized AI research environments. LangKit (WhyLabs, 2025) and NeMo Guardrails (Rebedea et al., 2023) also offer UQ scorers but are similarly narrow in scope. Lastly, while LM-Polygraph (Fadeeva et al., 2023) offers a robust collection of UQ-based approaches for LLMs, its documentation is written for a research audience and may lack ease of use for non-specialized practitioners.
We aim to bridge these gaps by introducing a comprehensive open-source Python package, uqlm, that democratizes advanced research in LLM uncertainty quantification. UQLM (Uncertainty Quantification for Language Models) implements a diverse array of uncertainty estimation techniques to compute generation-time, response-level confidence scores and uniquely integrates generation and evaluation processes. This integrated approach allows users to generate and assess content simultaneously, without the need for ground-truth data or external knowledge sources, and with minimal engineering effort. This democratization of access empowers smaller teams, researchers, and developers to incorporate robust hallucination detection into their applications, contributing to the development of safer and more reliable AI systems.
2 Usage
The uqlm library, available at https://github.com/cvs-health/uqlm, provides a collection of UQ-based scorers spanning four categories: black-box UQ, white-box UQ, LLM-as-a-Judge, and ensembles. [^2] For an detailed overview of available scorers and associated experiment results, we refer to the reader to this project’s companion paper, Bouchard and Chauhan (2025). The corresponding classes for these techniques are instantiated by passing an LLM object to the constructor. [^3] For the current version of uqlm, v0.1.8, a LangChain BaseChatModel is required. Note that an LLM is not required if users provide pre-generated responses and implement the score method. Each of these classes contains a generate_and_score method, which generates LLM responses to a user provided list of prompts and computes response-level confidence scores, which range from 0 to 1.
2.1 Black-Box Uncertainty Quantification
Black-box uncertainty quantification exploits the stochastic nature of LLMs and measures the consistency of multiple responses to the same prompt. These consistency measurements can be conducted with various approaches, including semantic entropy (Farquhar et al., 2024), non-contradiction probability (Chen and Mueller, 2024; Lin et al., 2024), BERTScore (Manakul et al., 2023; Zha et al., 2023), BLEURT (Sellam et al., 2020), exact match rate (Cole et al., 2023), and cosine similarity (Shorinwa et al., 2024). Black-box UQ scorers are compatible with any LLM, but increase latency and generation costs. The corresponding class for this collection of scorers is BlackBoxUQ.
Figure 1: Illustration of a Black-Box Scorer Workflow
To implement BlackBoxUQ.generate_and_score, users provide a list of prompts. For each prompt, an original response, along with additional candidate responses, are generated by the user-provided LLM, and consistency scores are computed using the specified scorers (see Figure 1). [^4] Note that Figure 1 depicts the approach for all black-box UQ scorers except semantic entropy, which does not designate an ‘original response’. If users set use_best=True, the uncertainty-minimized response is selected. [^5] Uncertainty-minimized response selection is based on semantic entropy (Farquhar et al., 2024). Below is a minimal example illustrating usage of BlackBoxUQ.
from uqlm import BlackBoxUQ
bbuq = BlackBoxUQ(llm=llm, scorers=["exact_match", "noncontradiction"])
results = await bbuq.generate_and_score(prompts=prompts, num_responses=5, use_best=True)
2.2 White-Box Uncertainty Quantification
White-box uncertainty quantification leverages token probabilities to compute uncertainty, as depicted in Figure 2. These approaches have the advantage of using the token probabilities associated with the generated response, meaning they do not add any latency or generation cost. However, because token probabilities are not accessible from all APIs, white-box scorers may not be compatible with all LLM applications. This collection of scorers can be implemented with the WhiteBoxUQ class, which offers two scorers: minimum token probability (Manakul et al., 2023) and length-normalized token probability (Malinin and Gales, 2021). Below is a minimal example of WhiteBoxUQ usage.
from uqlm import WhiteBoxUQ
wbuq = WhiteBoxUQ(llm=llm, scorers=["min_probability"])
results = await wbuq.generate_and_score(prompts=prompts)
Figure 2: Illustration of a White-Box Scorer Workflow
2.3 LLM-as-a-Judge
LLM-as-a-Judge uses an LLM to evaluate the correctness of a response to a particular question. To achieve this, a question-response concatenation is passed to one or more LLMs along with instructions to score the response’s correctness using the LLMPanel class (see Figure 3). In the constructor, users pass a list of LLM objects to the judges argument and specify one of four scoring templates for each judge with the scoring_templates argument. These four scoring templates are as follows: binary ($\{\text{incorrect},\text{correct}\}{ incorrect , correct }$ as $\{0,1\}{ 0 , 1 }$), ternary ($\{\text{incorrect},\text{uncertain},\text{correct}\}{ incorrect , uncertain , correct }$ as $\{0,0.5,1\}{ 0 , 0.5 , 1 }$), continuous (any value between 0 and 1), and a 5-point Likert scale (0,0.25,,100.2510,0.25,...,10 , 0.25 , … , 1). Implementing the generate_and_score method returns the score from each judge and aggregations of these scores, including minimum, maximum, average, and median. See below for a minimal example.
Figure 3: Illustration of LLM-as-a-Judge Workflow
from uqlm import LLMPanel
panel = LLMPanel(llm=llm1, judges=[llm2, llm3], scoring_templates=["continuous", "likert"])
results = await panel.generate_and_score(prompts=prompts)
2.4 Ensemble Approach
Lastly, uqlm offers both tunable and off-the-shelf ensembles that leverage a weighted average of any combination of black-box UQ, white-box UQ, and LLM-as-a-Judge scorers. Similar to the aforementioned classes, UQEnsemble enables simultaneous generation and scoring with a generate_and_score method. Using the specified scorers, the ensemble score is computed as a weighted average of the individual confidence scores, where weights may be default weights, user-specified, or tuned (refer to Appendix A for tuning details). If no scorers are specified, the off-the-shelf implementation follows an ensemble of exact match, non-contradiction probability, and self-judge proposed by Chen and Mueller (2024).
3 Conclusions
This paper introduced uqlm, an open-source Python toolkit offering a collection of state-of-the-art uncertainty quantification techniques for LLM hallucination detection. We believe uqlm democratizes uncertainty quantification techniques from the literature, empowering practitioners to effectively detect hallucinations at generation time with minimal engineering effort. To get started with uqlm, we refer readers to the Github repository and Documentation site.
Author Contributions
Dylan Bouchard was the principal developer and researcher of the UQLM project, responsible for conceptualization, methodology, and software development of the UQLM library. Mohit Singh Chauhan helped lead research and software development efforts. David Skarbrevik, Ho-Kyeong Ra, Viren Bajaj, and Zeya Ahmad contributed to software development.
Conflict of Interest
The authors are employed and receive stock and equity from CVS Health® Corporation.
Acknowledgments
We wish to thank Piero Ferrante, Blake Aber, Matthew Churgin, Erik Widman, Robert Enzmann, and Xue (Crystal) Gu for their helpful suggestions.
Appendix A Ensemble Tuning.
In order to tune the ensemble weights prior to using the generate_and_score method, users must provide a list of prompts and corresponding ideal responses to serve as an ‘answer key’. The LLM’s responses to the prompts are graded with a grader function that compares against the provided ideal responses. If a grader function is not provided by the user, the default grader function that leverages vectara/hallucination_evaluation_model is used.
Figure 4: Illustration of Ensemble Tuning
Once the binary grades (‘correct’ or ‘incorrect’) are obtained, an optimization routine solves for the optimal weights according to a specified classification objective. The objective function may be threshold-agnostic, such as ROC-AUC, or threshold-dependent, such as F1-score. After completing the optimization routine, the optimized weights are stored as class attributes to be used for subsequent scoring. Below is a minimal example illustrating this process.
from uqlm import UQEnsemble
## ---Option 1: Off-the-Shelf Ensemble (Chen & Mueller, 2023)---
# uqe = UQEnsemble(llm=llm)
# results = await uqe.generate_and_score(prompts=prompts, num_responses=5)
## ---Option 2: Tuned Ensemble---
scorers = [ # specify which scorers to include
"exact_match", "noncontradiction", # black-box scorers
"min_probability", # white-box scorer
llm # use same LLM as a judge
uqe = UQEnsemble(llm=llm, scorers=scorers)
# Tune on tuning prompts with provided ground truth answers
tune_results = await uqe.tune(
prompts=tuning_prompts, ground_truth_answers=ground_truth_answers
# ensemble is now tuned - generate responses on new prompts
results = await uqe.generate_and_score(prompts=prompts)
results.to_df()